OpenStudy (anonymous):

How to find the magnitude for vector WX for W(8, 3, 2) and X(3, 4, 9)?

OpenStudy (tkhunny):

Have you considered subtracting the components?

OpenStudy (anonymous):

What do you mean by subtracting? Like subtracting the components from each other?

OpenStudy (anonymous):

If so, would I do 8 - 3 = 5, 3 - 4 = -1, and 2 - 9 = -7?

OpenStudy (tkhunny):

Yes. With absolute values, you have now defined the three edge lengths of a rectangular prism. You need only to calculate its diagonal. The Extended Pythagorean Theorem should do.